OpenStudy (anonymous):

Factories A and B produce computers. Factory A produces 4 times as many computers as factory B. The probability that an item produced by factory A is defective is 0.014 and the probability that an item produced by factory B is defective is 0.031. A computer is selected at random and it is found to be defective. What is the probability it came from factory A?

OpenStudy (kropot72):

\[P(came\ from\ factory\ A)=\frac{4\times 0.014}{(4\times 0.014)+0.031}=you\ can\ calculate\]